OpenStudy (anonymous):

how do you simplify 2 square root 3?

OpenStudy (anonymous):

2/v3 * v3/v3 (v3/v3 is 1) you will get 2v3/3

OpenStudy (shaik0124):

2 sqrt root 3 2*1.73

OpenStudy (anonymous):

well im trying to convert it to inches?

OpenStudy (jdoe0001):

you'd want to get rid of the radical in the denominator to simplify the rational you'd do that by multiply both sides, up and down, by something that squares the denominator, thus giving you a non-radical value \(\bf \cfrac{2}{\sqrt{3}} \implies \cfrac{2}{\sqrt{3}} \times \cfrac{\sqrt{3}}{\sqrt{3}}\)

OpenStudy (jdoe0001):

ohh wait.. shoot, I misread :P, you don't have a rational

OpenStudy (jdoe0001):

\(\bf 2\sqrt{3} \) to inches from what?

OpenStudy (anonymous):

no no I think I got it. thankyou guys!